OpenStudy (anonymous):

For each inequality, choose the statement that describes its solution. If applicable, give the solution. -2(v+7)+38(greater then or equal to sign) 4 (2-v) no solution v(greater then or equal to sign)= v(less then or equal to sign)= all real numbers are solutions

OpenStudy (anonymous):

help :(

OpenStudy (anonymous):

Shouldn't there be numbers after your second and third choices?

OpenStudy (anonymous):

nope, whatever answer you get you have to put it in the second or third box depending on the sign, it doensnt give the answer

OpenStudy (anonymous):

-2(v+7)+38 >= 4 (2-v) -2v-14+38 >= 8-4v -2(v+7)+38 >= 4 (2-v) -2v-14+38 >= 8-4v -2v+24 >= 8-4v v-12 <= -4+2v 12 <= -4v y>= -3 not sure long time didn't solve problem like this

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

any time :)